# MADNESS Deblender

### v1

@@ 2.2 Test data
The test dataset, hereafter TestFields, was also generated using BTK but excluding the galaxies that were used for training and validation. Unlike the IsolatedSims and BlendedSims datasets; which contain small postage stamps, here we simulate multiple galaxies in random positions over a much larger field and test the results of our deblending algorithm using the true centers of the galaxies. To simplify the pipeline, we leave an empty margin in the edges of the field so that we can extract $45\times 45$ pixel cutouts from the field around each galaxy in the field. We simulate fields of size $205\times 205$ pixels with the centers of galaxies placed randomly only in the inner $150\times 150$ pixels ($0.25$ arcmin2) and use rejection sampling to ensure that the centers of galaxies do not lie within the same pixel. The TestFields simulations contain $12$ to $20$ galaxies in each field, corresponding to $48$ to $80$ galaxies per arcmin2. The density of galaxies is higher than the typical values we expect for LSST so that the deblenders can be evaluated and compared to each other for a higher blending rate; tests repeated with a more LSST-like density showed consistent results. In total, we generate $6000$ such fields with a total of 95 9729597295\,97295 972 galaxies.
@@ 3 Method
MADNESS uses a combination of two NN architectures: a VAE and a normalizing flow (NF). The goal of the VAE is to learn lower-dimension latent space representations of galaxies so that it can be used as a generative model, while the NF is used to model the underlying aggregate posterior distribution of galaxies in the latent space. Once trained, we can sample from the distribution learned by the NF to obtain latent space representation of new galaxies and feed these low dimensional representations to the decoder of the VAE to simulate new images. Lanusse et al. (2021) showed that such an architecture can be used to simulate complex and realistic models of galaxies with a hierarchical Bayesian framework to embed physical information within deep learning models. Although the authors explicitly account for both the PSF and the noise, in our work we do not decouple the PSF. Instead, we take a more conservative approach and train and evaluate our model with a fixed PSF to increase the effect of blending and make the task more challenging for the deblender. Our goal is to obtain a MAP estimate just like Lanusse et al. (2019) but instead of modeling the prior at pixel level with an auto-regressive network, we perform dimensionality reduction with a VAE and model our data-driven prior in the latent space using an NF network that provides us with an explicit likelihood. In other words, we obtain the MAP solution in the lower-dimension latent space instead of the high-dimensional pixel space.
In this section, we explain the data preprocessing, architectures, and training of VAE and NF, and discuss the methodology of our deblender.
@@ 3.1 Data normalization
Simulated galaxies typically have a wide range of flux values in each pixel. Training NN directly with this data can make the training very unstable so we need a mechanism to normalize the data. Although Arcelin et al. (2020) have found that a non-linear normalization is ideal for deblending with VAEs, the loss function in such normalized space is difficult to interpret. Since we want to obtain the final MAP solution in the flux space, we chose a linear normalization for the entire dataset such that the range of flux values falls mostly between 0 and 1. To be more precise, we divide the flux in all pixels by a factor of $10^{4}$, which is sufficient to bring the range of pixel values to a level that can be handled by the network.
3.2 Variational Auto-encoders
Figure 3: VAE architecture. The conv1, conv2, conv3, and conv4 layers refer to the four convolutional layers of the encoder all of which use a kernel size of 5 and stride 2. Following the convolutional layers are the dense layers d1 and d2 that predict the parameters of the latent space distribution. The probabilistic layer p1 uses the output of layer d2 to construct the latent space $z$, which is a multivariate normal distribution of 16 dimensions. Finally, the decoder takes as input a sample from the latent space distribution and applies a set of dense layers d3 and d4, followed by transposed convolutional layers (convT1, convT2, convT3, and convT4) also with kernel size 5 and stride 2; Finally, the output of convT4 is cropped to match the size of the input.
@@ 3.2 Variational Auto-encoders
An autoencoder falls under a class of neural networks that learn an encoding of high-dimensional data in a lower-dimensional latent space. The architecture can be divided into three distinct parts: the encoder, the latent space, and the decoder. The traditional auto-encoder provides a deterministic value while projecting data into the latent space. The loss function focuses only on the reconstruction and does not include any constraints on the properties of the latent space. Since the latent space is unconstrained, it is difficult to use the decoder as a generative model because we have no information about the underlying latent space distribution to sample from. First introduced by Kingma & Welling (2013), VAE overcomes this limitation by predicting a distribution in the latent space instead of a deterministic output and imposing a regularization over this predicted distribution.
More formally, if $x$ represents the observed data, and $z$ denotes the hidden latent variables, we want to find the mapping from the latent space to data that minimizes the marginal likelihood. In a VAE, the decoder represents this mapping $p_{\theta}(x|z)$, which is parameterized by $\theta$, the weights of the network. The training involves finding the parameters $\theta$ that maximize the marginal likelihood of the data, given a prior $p(z)$, to obtain $\hat{\theta}$ as
@@
The integral in Equation 1 is usually not tractable, especially for high dimensions of $z$. To get around this problem, a second parametric function $q_{\phi}(z|x)$ is introduced that approximates the true posterior $p_{\theta}(z|x)$. In the context of the VAE, this is referred to as the encoder and is parameterized by the weights $\phi$. The encoder helps avoid expensive MCMC sampling and approximates the latent space posterior distribution with a single forward pass through the pre-trained network. This technique is known as amortized variational inference. We can rearrange the terms in the Kullback-Leibler (KL, Kullback & Leibler, 1951) divergence between this approximated posterior and the true posterior $p(z|x)$, to evaluate $\log p(x)$ or the evidence as
$\begin{split}\log p(x)=\operatorname{\mathbb{E}}_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\operatorname{\mathbb{D}}_{\rm{KL}}[q_{\phi}(z|x)||p(z)]\\ +\operatorname{\mathbb{D}}_{\rm{KL}}[q_{\phi}(z|x)||p(z|x)],\end{split}$ (2)
where the first term is the expectation of the likelihood and $\operatorname{\mathbb{D}}_{\rm{KL}}$ represents the KL divergence between the two distributions. Since the last term in Equation 2 is a KL divergence, it will always have a positive value. Equation 2 then leads to
$\log p(x)\geq\operatorname{\mathbb{E}}_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\operatorname{\mathbb{D}}_{\rm{KL}}[q_{\phi}(z|x)||p(z)].$ (3)
By using the trick of amortized variational inference, we have arrived at Equation 3 which is tractable and is called the evidence lower bound (ELBO). By maximizing the ELBO over all the observed data (in our case, the simulations) we can optimize values of parameters $\theta$ and $\phi$, i.e. the weights of the network.
In our implementation, we use the TensorFlow  [^9] https://www.tensorflow.org/ framework except for the probabilistic layers where we use TensorFlow-Probability  [^10] https://www.tensorflow.org/probability. As shown in Figure 3, the encoder has 4 convolutional layers with 64, 128, 256, and 128 filters. Each layer uses a convolution kernel of size 5 and a stride of 2. These layers are followed by 2 dense layers that predict the latent space distribution. Each layer in the encoder is followed by PReLU (He et al., 2015) activation functions except the final dense layer, which is left without an activation function as it predicts the parameters for a multivariate normal distribution layer that describes the $16$-dimensional latent space distribution. The latent space is followed by the decoder, which also has two dense layers and 4 transposed convolutional layers with 128, 96, 64, and 6 filters. The first 3 layers use a kernel of size 5 and stride 2 while the final layer has a kernel of size 3 and stride 1. Once again we use the PReLU activation function in all the layers except the final layer where we apply a ReLU (Nair & Hinton, 2010) activation function to prevent negative predictions of flux. Finally, the output is cropped to match the size of the input. In all, the encoder and decoder have around $6.7$M and $3.5\rm{M}$ trainable parameters, respectively. The architecture of the encoder is chosen to be more complex because it is later trained as a deblender to obtain the initial point for gradient descent, as described in Section 3.4.
Optimization is done by minimizing the ELBO loss function. However, as observed very frequently in VAE literature (Chen et al., 2017), the KL divergence term in ELBO dominates and prevents the network from properly reconstructing galaxies. Among several different approaches, typically annealing-based methods are used to slowly increase the weight of the KL divergence term during training. Alternatively, this dominance can also be mitigated by multiplying the KL term with another hyperparameter $\beta$ in the loss function, to adjust the relative weight between the reconstruction and the regularization terms. These models are called the $\beta$-VAE (Higgins et al., 2017). By using $\beta<1$ we can improve the reconstruction fidelity but this also implies that the learned aggregate posterior will differ from the prior. An alternative approach could be to remove the KL term entirely. However, the lack of regularization will lead to a more complex latent space distribution as there is no incentive for the VAE to follow the ELBO-prior. The added complexity of the latent space would make it more difficult to perform a gradient descent in the VAE latent space for the MAP estimate. After testing the $\beta-$VAE for several values of the hyperparameter $\beta$, we chose the value of $0.01$ as no significant improvement was observed by lowering the value further.
Even after fixing the weight of KL divergence, we observed that the reconstructions were still poor in the u-band. This is because the VAE starts learning features in the u-band only in later iterations due to the relatively lower SNR of galaxies compared to the other bands. To address this challenge, we include a warm-up phase in the VAE training with a slightly modified loss function. We introduce the structural similarity index measure (SSIM, Wang et al., 2004) in a way that the NN is forced to predict correct structural information in all the bands from the initial iterations. The SSIM is a metric typically used in image processing to assess the similarity between two images. An SSIM value close to $1$ represents highly correlated structural information between the two images, while $-1$ represents anti-correlation. Using SSIM we define the warm-up loss as
where $\operatorname{\mathbb{S}}$ denotes the average SSIM across filters, computed using input and output images, $\alpha$ is a weight applied to SSIM, and $\beta$, the KL divergence coefficient, is set to $0.01$. The coefficient $\alpha$ starts at $1$ and is decreased by $0.02$ in every epoch until it reaches a value of 00 after $50$ epochs. The SSIM is computed on normalized images where each band of input and output is divided by the maximum pixel value in that band so all values are in the range $[0,1]$. Since the $\operatorname{\mathbb{S}}$ is equally weighted among all the bands, the network is forced to learn features in all the filters from the very beginning.
We use the Adam optimizer (Kingma & Ba, 2017) with an initial learning rate of $10^{-5}$, along with a learning rate scheduler that decreases the learning rate by a factor of $2.5$ every $40$ epochs. Additionally, since we have only a linear normalization of our data and some galaxies are extremely bright compared to others, we introduce a clip value of $0.1$ on the gradients obtained during the optimization to stabilize the training of our network. After the initial warm-up phase is complete, we drop the SSIM term from the loss function and restart the training with a learning rate of $10^{-5}$ for a maximum of $200$ epochs with an early stopping patience of $20$ so that the network stops training if the loss has not decreased over $20$ iterations. Even during this training phase, we keep the same learning rate scheduler and clip value.
Once trained, the VAE can be used as a generative model by sampling from the latent space prior $p(z)$ in the ELBO, and feeding it to the decoder. However, this assumes that the prior is representative of the aggregate posterior which is not strictly valid for a VAE, and introducing a $\beta<1$ in the $\beta-$VAE makes the aggregate posterior depart further from the prior. Therefore, sampling from the prior would lead to biased and out-of-distribution samples. Engel et al. (2017) proposed using GANs to model the latent space to generate high-fidelity images GANs, however, do not provide a way to evaluate the likelihood necessary for the MAP estimate. As an alternative to GANs, in Section 3.3 we discuss in detail another generative model called normalizing flows (NFs) that gives us access to a tractable likelihood.
3.3 Normalizing Flows
Figure 4: Schematic of the training of the normalizing flow. The encoder is the pre-trained network illustrated in Figure 3 and the normalizing flow consists of 6 bijective transformations $f_{\varphi_{i}}$ where the $i^{th}$ transformation is parameterized by $\varphi_{i}$, which represents the weights of the MAF layer containing (32, 32) units with tanh activations followed by a permutation layer.
NFs are capable of modeling a complex distribution by sequentially applying bijective transformations to a simpler distribution that can easily be sampled. Mathematically, if we consider $\operatorname{\tilde{X}}$ and $\operatorname{\tilde{Z}}$ to be two random variables related by bijective mapping $f_{\varphi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, parameterized by $\varphi$, such that $\operatorname{\tilde{X}}$ = $f_{\varphi}(\operatorname{\tilde{Z}})$, we can relate the transformed distribution $p_{\operatorname{\tilde{X}}}(\operatorname{\tilde{x}};\varphi)$ to the prior $p_{\operatorname{\tilde{Z}}}(\operatorname{\tilde{z}})$, using the change of variable formula to compute the probability of sample $\operatorname{\tilde{x}}$ from the data as
$p_{\operatorname{\tilde{X}}}(\operatorname{\tilde{x}};\varphi)=p_{\operatorname{\tilde{Z}}}(f_{\varphi}^{-1}(\operatorname{\tilde{x}}))\left|\text{det}\left(\frac{\partial f_{\varphi}^{-1}(\operatorname{\tilde{x}})}{\partial\operatorname{\tilde{x}}}\right)\right|.$ (5)
The advantage of NFs is that they provide a tractable likelihood. The goal of fitting the normalizing flow is to find appropriate parameters $\varphi$ so that the transformed distribution is the same as the target distribution we are trying to model, which in our case is the aggregate posterior in the VAE latent space. The optimization can be done by minimizing the negative log-likelihood:
$\hat{\varphi}=\operatorname*{arg\,min}_{\varphi}\left[-\log p_{\operatorname{\tilde{X}}}(\operatorname{\tilde{x}};\varphi)\right];$ (6)
$\hat{\varphi}=\operatorname*{arg\,min}_{\varphi}\left[-\log p_{\operatorname{\tilde{Z}}}(f_{\varphi}^{-1}(\operatorname{\tilde{x}}))-\log\left|\text{det}\left(\frac{\partial f_{\varphi}^{-1}(\operatorname{\tilde{x}})}{\partial\operatorname{\tilde{x}}}\right)\right|\right]\,.$ (7)
Provided that the transformation $f_{\varphi}$ is flexible enough, we can model a complex data distribution using this minimization. In our work, we use a family of NFs called Masked AutoRegressive Flows (MAF) presented by Papamakarios et al. (2017). The MAF architecture uses autoregressive transformations in which
$\operatorname{\tilde{x}}_{i}=\operatorname{\tilde{z}}_{i}\sigma_{\theta_{i}}+\mu_{\theta_{i}},$ (8)
where $\sigma_{\theta_{i}}$ and $\mu_{\theta_{i}}$ are functions of $\operatorname{\tilde{x}}_{1}....\operatorname{\tilde{x}}_{i-1}$, parameterised by $\theta_{i}$ which is typically a dense network.
@@ 3.3 Normalizing Flows
Choosing a transformation like the MAF makes the computation of the determinant of the Jacobian tractable. But one layer of MAF is not expressive enough to model a non-trivial distribution and results depend heavily on the order of variables. To overcome these limitations, we must apply a set of such transformations to the base distribution, while performing permutations to the order of variables between layers, i.e.
$\operatorname{\tilde{x}}=f_{\varphi}(\operatorname{\tilde{z}})=f_{\varphi_{n}}...\circ f_{\varphi_{2}}\circ f_{\varphi_{1}}(\operatorname{\tilde{z}})\,,$ (9)
@@
where $f_{\varphi_{j}}$ represents the $j^{\text{th}}$ layer of transformation.
Figure 5: Comparison of the latent space distribution of 50 0005000050\,00050 000 samples each from the validation set of IsolatedSims and the distribution learned by the normalizing flow (NF). As we move out from the center of the distribution, the contours represent $25\%$, $50\%$, $75\%$, and $90\%$ of the probability mass.
We use TFP to create bijective autoregressive networks with 2 hidden layers containing 32 units each followed by $\tanh$ activation functions, to be used within MAF layers. We use 6 sets of alternating MAF and permutation layers, resulting in a model with a total of 15 9361593615\,93615 936 free parameters.
@@
We train the model by minimizing the negative log-likelihood of the latent space representations of galaxies in IsolatedSims obtained by a forward pass through the trained encoder as shown in Figure 4. We use the Adam optimizer and start with a learning rate of $10^{-4}$, clip value of $0.01$, and the same setting for the learning rate scheduler and patience as the VAE, for a maximum of $200$ epochs.
(a) Galaxies simulated with VAE after sampling from the ELBO prior.
(b) Galaxies simulated after sampling from the distribution learned using the normalizing flow.
@@
Figure 6: Examples of randomly chosen single galaxy simulations in the r-band using the trained decoder as the generative model. The color bar for each plot shows the normalized electron counts in the pixels.
@@
Thus, we train the NF network to model the target distribution, which in our case happens to be the aggregate posterior of galaxies in the latent space of the trained VAE. In Figure 5, we show that the trained normalizing flow successfully models the underlying latent space representation of the galaxies. Figure 6 shows the impact on simulating galaxies with the VAE when shifting to an NF instead of the ELBO prior. We can conclude from Figure 6 that sampling from the standard normal prior used to train the VAE fails to provide us with realistic-looking galaxies but sampling from the aggregate posterior as modeled by the NF, gives us more reliable simulations (Lanusse et al., 2021; Arcelin, 2021). We now use this architecture to obtain the log prior required for the MAP solution.

### v2

@@ 2.2 Test data
The test dataset, hereafter TestFields, was also generated using BTK but excluding the galaxies that were used for training and validation. Unlike the IsolatedSims and BlendedSims datasets; which contain small postage stamps, here we simulate multiple galaxies in random positions over a much larger field and test the results of our deblending algorithm using the true centers of the galaxies. We simulate fields of size $205\times 205$ pixels with the centers of galaxies placed randomly only in the inner $150\times 150$ pixels ($0.25$ arcmin2) to ensure that all pixels of the predicted galaxies lie within the field. We used rejection sampling to ensure that the centers of galaxies do not lie within the same pixel. The TestFields simulations contain $12$ to $20$ galaxies in each field, corresponding to $48$ to $80$ galaxies per arcmin2. The density of galaxies is higher than the typical values we expect for LSST so that the deblenders can be evaluated and compared to each other for a higher blending rate; tests repeated with a more LSST-like density showed consistent results. In total, we generate $6000$ such fields with a total of 95 97295\,972 galaxies.
@@ 3 Method
MADNESS uses a combination of two neural network architectures: a VAE and a normalizing flow. The goal of the VAE is to learn low-dimension latent-space representations of galaxies so that it can be used as a generative model, while the normalizing flow is used to model the underlying aggregate posterior distribution of galaxies in the latent space. Once trained, we can sample from the distribution learned by the normalizing flow to obtain a latent space representation of new galaxies and feed these low-dimensional representations to the decoder of the VAE to simulate new images. Lanusse et al. (2021) showed that such an architecture can be used to simulate complex and realistic models of galaxies with a hierarchical Bayesian framework to embed physical information within deep learning models. Although the authors explicitly account for both the PSF and the noise, in our work, we do not decouple the PSF. Instead, we take a more conservative approach and train and evaluate our model with a fixed PSF to increase the effect of blending and make the task more challenging for the deblender. Our goal is to obtain a MAP estimate similar to that of Lanusse et al. (2019) but instead of modeling the prior at pixel level with an auto-regressive network, we perform dimensionality reduction with a VAE and model our data-driven prior in the latent space using a normalizing flow network that provides with an explicit likelihood. In other words, we obtain the MAP solution in the low-dimensional latent space instead of the high-dimensional pixel space. In this section, we explain the data preprocessing, architectures, and training of VAE and normalizing flow, and discuss the methodology of our deblender.
@@ 3.1 Data normalization
Simulated galaxies typically have a wide range of flux values in each pixel. Training the neural network directly with this data can make the training very unstable, so we need a mechanism to normalize the data. Although Arcelin et al. (2020) have found that a non-linear normalization is ideal for deblending with VAEs, the loss function in such a normalized space is difficult to interpret. Since we want to obtain the final MAP solution in the flux space, we choose a linear normalization for the entire dataset such that the range of flux values falls mostly between 0 and 1. To be more precise, we divide the flux in all pixels by a factor of $10^{4}$, which is sufficient to bring the range of pixel values to a level that can be handled by the network. Studies with other normalization techniques are left as future work.
3.2 Variational autoencoders
@@
An autoencoder falls under a class of neural networks that learn an encoding of high-dimensional data in a low-dimensional latent space. The architecture can be divided into three parts: the encoder, the latent space, and the decoder. The traditional auto-encoder provides a deterministic value while projecting data into the latent space. The loss function focuses only on the reconstruction and does not include any constraints on the properties of the latent space. Since the latent space is unconstrained, it is difficult to use the decoder as a generative model because we have no information about the underlying latent space distribution to sample from. First introduced by Kingma & Welling (2013), VAE overcomes this limitation by predicting a distribution in the latent space instead of a deterministic output and imposing a regularization over this predicted distribution.
More formally, if $x$ represents the observed data and $z$ denotes the hidden latent variables, the decoder is represented by the mapping $p_{\theta}(x|z)$, while the encoder is represented by a second parametric function $q_{\phi}(z|x)$, where $\theta$ and $\phi$ refer to the weights of the network.
The optimization step of the VAE involves finding the parameters $\hat{\theta}$ and $\hat{\phi}$ that minimize the evidence lower bound (ELBO) on the right-hand side of the inequality
$\log p(x)\geq\operatorname{\mathbb{E}}_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\operatorname{\mathbb{D}}_{\rm{KL}}[q_{\phi}(z|x)||p(z)],$ (1)
where $\log p(x)$ is called the evidence, the first term is the expectation of the likelihood, $p(z)$ is the latent space prior, and $\operatorname{\mathbb{D}}_{\rm{KL}}$ represents the Kullback–Leibler (KL) divergence which measures how different the two distributions are.
Optimization is done by minimizing the ELBO loss function. However, as observed very frequently in VAE literature (Chen et al., 2016), the KL divergence term in ELBO dominates and prevents the network from properly reconstructing galaxies. Among several different approaches, typically annealing-based methods are used to slowly increase the weight of the KL divergence term during training. Alternatively, this dominance can also be mitigated by multiplying the KL term with another hyperparameter $\beta$ in the loss function, to adjust the relative weight between the reconstruction and the regularization terms. These models are called the $\beta$-VAE (Higgins et al., 2017).
After lowering the weight of KL divergence, we observed that the reconstructions were still poor in the u-band. This is because the VAE starts learning features in the u-band only in later iterations due to the relatively lower S/N of galaxies compared to the other bands. To address this challenge, we include a warm-up phase in the VAE training with a slightly modified loss function. We introduce the structural similarity index measure (SSIM, Wang et al., 2004) in a way that the neural network is forced to predict correct structural information in all the bands from the initial iterations. The SSIM is a metric typically used in image processing to assess the similarity between two images. An SSIM value close to $1$ represents highly correlated structural information between the two images, while 1-1 represents anti-correlation. Using SSIM, we define the warm-up loss as
where $\operatorname{\mathbb{S}}$ denotes the average SSIM across filters, computed using input and output images of each band normalized to a maximum value of 1, and $\alpha$ is a weight applied to SSIM. In practice, we set $\alpha$ in a way that it reduces to zero after a few initial iterations, as described in Appendix B.1. The exact architecture of the VAE and hyperparameters of the model used to obtain our results is presented in Appendix B.1.
Once trained, the VAE can be used as a generative model by sampling from the latent space prior $p(z)$ in the ELBO, and feeding it to the decoder. However, this assumes that the prior is representative of the aggregate posterior which is not strictly valid for a VAE, and introducing a $\beta<1$ in the $\beta$-VAE makes the aggregate posterior depart further from the prior. Therefore, sampling from the prior would lead to biased and out-of-distribution samples. Engel et al. (2017) proposed using GANs to model the latent space to generate high-fidelity images. GANs, however, do not provide a way to evaluate the likelihood necessary for the MAP estimate. As an alternative to GANs, in Section 3.3 we discuss in detail another generative model called normalizing flows that gives access to a tractable likelihood.
3.3 Normalizing flows
@@
In our work, we used a family of normalizing flows, called masked autoregressive flows (MAFs), presented by Papamakarios et al. (2017), which are typically parameterized by a dense network. However, one layer of MAF is not expressive enough to model a non-trivial distribution, and results depend heavily on the order of variables. Thus, we must apply a set of such transformations to the base distribution, while performing permutations to the order of variables between layers, i.e.
[…]
@@
where $f_{\varphi_{j}}$ represents the $j^{\text{th}}$ layer of transformation, parameterized by a neural network $\varphi_{j}$.
@@
We train the model by minimizing the negative log-likelihood of the latent space representations of galaxies in IsolatedSims obtained by a forward pass through the trained encoder as shown in Figure 15. Appendix B.2 describes the network architecture and training hyperparameters for the model used to generate the results hereafter.
Figure 3: Comparison of the latent space distribution of 50 00050\,000 samples each from the validation set of IsolatedSims and the distribution learned by the normalizing flow. As one moves out from the center of the distribution, the contours represent $25\%$, $50\%$, $75\%$, and $90\%$ of the probability mass.
@@
Figure 4: Examples of randomly chosen single galaxy simulations in the r-band using the trained decoder as the generative model. Panel (a) shows galaxies simulated with VAE after sampling from the ELBO prior, while panel (b) shows galaxies simulated after sampling from the distribution learned using the normalizing flow. The color bar for each plot shows the normalized electron counts in the pixels.
@@
In Figure 3, we show that the trained normalizing flow successfully models the underlying latent space representation of the galaxies. Figure 4 shows the impact on the simulation of galaxies with the VAE when shifting to a normalizing flow instead of the ELBO prior. We can conclude from Figure 4 that sampling from the standard normal prior used to train the VAE fails to generate realistic-looking galaxies, but sampling from the aggregate posterior as modeled by the normalizing flow gives more reliable simulations (Lanusse et al., 2021; Arcelin, 2021). We now use this architecture to obtain the log prior required for the MAP solution.